The Bakers at Healthy Bakery can make 280 bagels in 5 hours. How many bagels can they bake in 14 hours? What is that rate per hour?
1
Expert's answer
2019-05-20T03:23:30-0400

Bakers can make 280 bagels in 5 hours so basically their rate per hour is

"n = 280\/5 = 56 bagels\/ hour"

In accordance to this rate in 14 hours they will produce

"N=14*n=14*56=784 bagels"